A Stirling-engine is used in the heat-pump mode to heat a house. The engine maintains a temperature of 20.5 °C inside the house. The temperature of the Earth loop is 14.1 °C. (The Earth loop buried deep under the ground is the cold reservoir of this heat pump.) What is the coefficient of performance of this heat pump?

Tries 0/12

If the power of the electric motor driving the heat pump is 157 W, then what is the rate at which heat is delivered to the house?
Answer #1

(1 cop of stirling Engine. Ti = 20.5°C T2= 14,1°C = (2005+273) K = 293.5K =(141+273 ) K = 287.1 K Mow COP (Heat Remp) = TH =

Please rate
if any mistake in this answer please comment i will clarify your doubt . thank you